OpenStudy (anonymous):

hat is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)?

OpenStudy (anonymous):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)? x2 + y2 − 4x + 2y + 1 = 0 x2 + y2 + 4x − 2y + 1 = 0 x2 + y2 + 4x − 2y + 9 = 0 x2 − y2 + 2x + y + 1 = 0

OpenStudy (anteater):

(x - h)^2 + (y - k)^2 = r^2 ; (h,k) = (-2,1) and r is the distance between (-2,1) and (-4,1) So then plug in the values given, expand the squared expressions and subtract r^2 from both sides of the equation to get your equation in general form.

OpenStudy (anteater):

Distance between (-2,1) and (-4,1) is 2. (x - (-2))^2 + (y - 1)^2 = (2)^2 --> (x+2)^2 + (y-1)^2 = 4 --> x^2 + 4x + 4 + y^2 - 2y + 1 = 4 --> x^2 + 4x + y^2 - 2y + 5 = 4 --> x^2 + 4x + y^2 - 2y + 1 = 0

OpenStudy (anteater):

Or x^2 + y^2 + 4x - 2y + 1 = 0